OpenStudy (anonymous):

let's say we have a function F(y(t),t)=y*G(t) the partial derevative of F with respect to t is y*dG/dt or dy/dt*G+dG/dt*y I would really apreciate an explanation

OpenStudy (anonymous):

option 2

OpenStudy (anonymous):

since both are function of t!

OpenStudy (abb0t):

\[F(y(t), t) = y \times G(t)\] ?

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

\[\frac{ ∂G }{ ∂t }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if both G and Y are function of t.. then you have to use the product rule!

OpenStudy (anonymous):

mashy our teacher used the first one in determining the solution of a differential equation of first order that's what makes me crazy

OpenStudy (anonymous):

no he mentioned its a function of time see the question!!

OpenStudy (anonymous):

y is a function of time

OpenStudy (anonymous):

Then Y must be independent of t... !!

OpenStudy (abb0t):

stop yelling @ me!! Ur hurting my feelings, bro :'(

OpenStudy (anonymous):

awww.. come here you !!

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

wow i get a medol and abbot gets none.. yay yay :P

OpenStudy (anonymous):

who the hell gave u a medal i think your answer is wrong man here we're talking about partial derevative which is diffrent from derevating everything with respect to t

OpenStudy (anonymous):

Is it the partial derivative of \[ F(u, t) \]Or of \[ F(y(t),t) \]?

OpenStudy (anonymous):

the second one bro partial derevative with respect to t

OpenStudy (anonymous):

Well it was ambiguous, and \(F(u, t) \neq F(y(t),t)\)

OpenStudy (anonymous):

and so ?

OpenStudy (anonymous):

if both functions are time dependent.. then really there is no difference between partial and normal derivatives.. HELL why would you even DO a partial derivative doesn't even make any sense :-/

OpenStudy (anonymous):

-_- it makes sense when u'r solving a diferential equation of first order first you find F(y,t) then you derevative partially with respect to y or t and force it to be equal to the equation you have left so yeh u need a partial derevative

OpenStudy (anonymous):

give me a differential equation!!

OpenStudy (anonymous):

Like @Mashy is saying, \(F(y(t),t)\) doesn't have a partial derivative because it is not a multiple variable function. It's the OUTPUT of \(F(u, t)\) which happens to be a single variable function.

OpenStudy (abb0t):

well, that escalated quickly.

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

M(y,t)dt+N(y,t)dy=0 where M=df/dt and N=df/dy

OpenStudy (abb0t):

Whoa, now we're doing ODE's?

OpenStudy (anonymous):

thats called as EXACT FORM right?

OpenStudy (abb0t):

Correct, Mashy :)

OpenStudy (anonymous):

wio i didnt catch up with what you said i mean here we got 2 variable well not two variables y is a function of time and hell

OpenStudy (anonymous):

well then in that.. case.. i have forgetting how to do it :D

OpenStudy (abb0t):

This is escalating too damn quickly.

OpenStudy (anonymous):

it's like i have holes in mathematics so when i advance i fall into some of them

OpenStudy (anonymous):

holes in mathematics.. lol funny :D

OpenStudy (anonymous):

Can you give us \(M, N\)?

OpenStudy (anonymous):

he already mentioned.. df/dy and df/dt!

OpenStudy (abb0t):

Well, you want to find the PD for either M or N first. I think it's ∂M/∂x and ∂N/∂y ?

OpenStudy (anonymous):

it doesn't matter what M and N are the only restriction is that M and N are the partial derevatives of a certain function F(y,t)

OpenStudy (anonymous):

wow i did this 4 years back!! ..

OpenStudy (anonymous):

so i don't remember much :-/

OpenStudy (anonymous):

sadly

OpenStudy (abb0t):

Well, you have to check that they are exact first for this to work.

OpenStudy (abb0t):

Otherwise, you are wasting time trying to do furhter work than is necessary.

OpenStudy (anonymous):

\[ \begin{array}{rcl} M(y,t)dt+N(y,t)dy &=& 0 \\ M(y,t)dt &=& -N(y,t)dy \end{array} \] I would try integrating \(M\) with respect to \(t\) and then differentiating with respect to \(y\).

OpenStudy (anonymous):

wait can i post a picture ?

OpenStudy (anonymous):

yup i can wait

OpenStudy (anonymous):

really?!? of yourself?? we are in the middle of math here!

OpenStudy (anonymous):

here check tht

OpenStudy (anonymous):

i'm too handsome :p

OpenStudy (anonymous):

so i have heard :P

OpenStudy (anonymous):

i mean .. i did.. just now :P.. from you!!

OpenStudy (anonymous):

so any idea about what i posted?

OpenStudy (abb0t):

What it looks like you're trying to do in that step is find a integrating factor to make the solution exact?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

wait wait wait we found the integrating factor we're in the middle of finding F

OpenStudy (anonymous):

see where's the arrow pointing

OpenStudy (abb0t):

I'm not quite sure what you did there, I'm used to doing it a different method..

OpenStudy (anonymous):

tht's the teacher's doing not me :3

OpenStudy (anonymous):

oh well i'm off now i'll check again after half an hour or so thanks ppl

OpenStudy (anonymous):

i'm back people any news,